### v1

# Mesoscopic Modeling of High-Density Carbon Nanotube Films for Memristive Device Applications

Abstract
Carbon nanotube (CNTs) materials, which exhibit intrinsically high electrical conductivity, are promising candidates for energy-efficient electronic devices. Recently, high-density CNT films have also been successfully employed as switching elements in non-volatile memory cells. However, the mechanism of electrical conduction through such complex systems is still poorly understood. To identify structural parameters that govern the electrical current in CNT films, we employed coarse-grained molecular dynamics to construct dense mesoscale CNT film models, where we considered CNTs with different chiralities and lengths. The effects of CNT geometrical features on the film morphologies were quantified by devising a set of structural descriptors and analyzing their mutual correlations. The impact of varying the concentration of amorphous carbon (aC) inclusions on the film structure was assessed. Finally, we employed a nodal analysis framework to compute the electrical current across the networks and correlate the charge transport characteristics to the underlying structural descriptors. Transport is found to be enhanced in films that exhibit high curvature and buckling, low bundling, and strong connectivity, with amorphous carbon components playing a nontrivial configuration-dependent role. These findings provide a framework for the rational design of CNT-based memristor architectures and highlight the potential of mesoscale modeling to guide the engineering of advanced nanostructured materials.
keywords:
High-Density Carbon Nanotube Films, Memristive Devices, Amorphous Carbon, Mesoscopic Modeling, Electrical Current, Structural Descriptors, Morphology-Transport Correlation1 Introduction
Carbon-based electronics is widely regarded as a promising alternative to conventional silicon-based devices 1, 2, 3, 4, 5, 6, 7. In particular, carbon materials such as carbon nanotubes (CNTs) and graphene exhibit exceptional electronic, optoelectronic and mechanical properties 8, 9, 10. The extremely high predicted tensile strength and Young’s modulus of CNTs make them among the strongest and stiffest known materials 11. Although the properties of single and small ensembles of CNTs have been extensively studied in recent decades, both experimentally and theoretically, the properties of CNT films, which are strongly affected by the formation of bundles and junctions 2, 12, 13, 14, 15, 16, remains comparatively less understood 17, 18.
Experimentally, it has been shown that the properties of CNT films can strongly depend on their density 19, 20, 21, 22, 23, 24. Various strategies have been developed to increase the density of CNT films, such as annealing to increase ordering and crystallinity 19, 21 or mechanical pressure to promote tighter packing 23. For instance, buckypapers (i.e. macroscopic aggregates of carbon nanotubes) exhibit almost linearly improved electrical and thermal conductivity under compression 23. The fabrication method also seems important, as sequential layering of CNT films can be used to tune the properties of the final film 25, 9. Finally, although less frequently discussed, amorphous carbon is expected to be inherently present between tubes and bundles as a byproduct of the CNT growth process, but its effect on the physical properties of CNT films remains poorly understood.
Recently, CNT films have found new applications in resistive random-access memory (RRAM) devices, which are considered promising candidates for the next generation of nonvolatile memory (NRAM) technologies 26, 27. In such devices, the memory state information is encoded in the conductivity of the NRAM cell determined by reversible changes in cell resistance. Resistive switching between these states is governed by a complex interplay between local structural features and electron transport pathways at the nanoscale. Understanding the mechanisms of electron transport through CNT films remains challenging due to their intrinsically disordered structure, heterogeneous connectivity, and possible presence of amorphous carbon.
Mesoscale modeling offers an attractive framework to address these issues 28, 29, 30, 31. It allows the direct incorporation of structural complexity, such as bundling, curvature, and intertube junctions into current-carrying networks, without relying on computationally expensive atomistic simulations or overly simplified continuum models. In particular, the use of structural descriptors may provide a compact and interpretable representation of key morphological features, thereby enabling a systematic exploration of how geometry and topology influence electron transport.
Early theoretical work has highlighted the importance of geometric morphology in determining the physical properties of CNT films 32, 33. However, these studies focused on two-dimensional (2D) networks and did not include an atomistic description of the film structure. More recently, computational modeling has been used to understand the behavior of CNT films at a mesoscopic level. The methodology can generally be divided into two classes: (i) the use of a mesoscopic tubular (MT) potential that describes inter-CNT interactions based on a continuous interatomic potential 28, 34, 30, 35, and (ii) the mesoscopic distinct element method (MDEM), which enables the simulation of interactions between particles of arbitrary shapes 36, 37, 38, 39. Simulations using these mesoscopic models have shown that various factors, particularly those related to underlying structure of the network, have a strong influence on mechanical and thermal properties 34, 40, 31, 35, 39. In contrast to the 2D approximation representing CNTs as straight lines, the MT potential allows one to directly study the effects of CNT bending and bundling, as well as the construction of more complex structures 29, 30, 34, 41, 42, 43. For example, it has been shown that increased bending energy leads to an overall increase in film stability, defined as the tendency not to disintegrate into cellular structures consisting of individual bundles weakly connected to each other 34.
In this study, we construct and analyze high-density CNT films with varying chirality, length, density, and aC content to investigate whether such films can be rationally designed to achieve specific target morphologies. Unlike previous studies that focused on relatively low-density systems, we consider films with density approaching the close-packed limit and study both single-layer and stacked multi-layer configurations, thereby allowing us to assess the influence of the fabrication method. We employ a mesoscale modeling approach capable of capturing mechanical deformation and local disorder. By quantifying each structure through a set of physically motivated descriptors, we establish correlations between structural properties and electron transport efficiency. Moreover, we develop a simple approach to mimic the effect of amorphous carbon inside the films. Overall, this work sheds light on the structural factors governing conductivity in complex CNT networks and provides a foundation for future modeling of resistive switching mechanisms in CNT-based memristive devices.
The paper is organized as follows. We discuss the choice of structural parameters, such as tube length and film density, in Section 2.1. The mesoscopic model and the incorporation of aC into the simulations are presented in Section 2.2. The procedure for constructing the films is presented in Section 2.3. We define the set of descriptors in Section 2.4. The nodal analysis used to compute the electrical current through the films is detailed in Section 2.5. The results of the simulations are presented in Section 3 followed by concluding remarks in Section 4. Additional data and supporting figures are provided in the Supplemental Information (SI).
2 Methodology
2.1 Structural parameters
The CNT films targeted in this study are modeled after experimental systems fabricated as described in Refs 27, 44. Due to the specifics of the growth and deposition techniques, the formed CNT films consist of stacked layers with different CNT lengths and chirality distributions (see SI). Based on the experimentally measured distributions, we selected three representative lengths: a short one (15 nm), a intermediate one (40 nm), and a long one (100 nm, corresponding to half the size of the simulation box), as well as two different chiralities, (16,0) and (32,0), each associated with a specific diameter: and .
Experimentally fabricated films are predominantly composed of multi-walled CNTs, whereas simulated films are composed of single-walled CNTs. We chose a high density of 0.6 gcm-3 (the close-packed density for (32,0) tubes is 0.8 gcm-3), and also simulate films at the lower density of 0.3 gcm-3 for comparison. The films were constructed either as single-layer or as four-layer stacked configurations. Finally, the amorphous carbon content was experimentally estimated to lie between 10 and 40% by mass. Starting from two reference configurations, namely 100 nm CNT length, single-layer, 0.6 gcm-3, and either (16,0) or (32,0), the amount of aC is varied from 0% to 40% with a 10% increment. By keeping constant the target density, the number of CNTs in the simulation box decreases as the amount of aC particles increases. The standard deviation of the final density is 710-4 gcm-3 for the 0.3 gcm-3 samples and 410-4 gcm-3 for the 0.6 gcm-3 samples. Altogether, a total of 32 distinct structures were constructed and analyzed.
2.2 Mesoscopic Model
CNTs are modeled as collections of point particles (beads), each representing a 1.0 nm-long cylindrical segment of a CNT; the terminal beads are lighter and only represent a 0.5-nm segment. Beads within the same CNT are connected by harmonic stretching and bending potentials, while beads from different, randomly oriented CNTs interact through nonbonded Lennard-Jones potentials, according to the mesoscopic tubular force field developed by Volkov, Zhigilei, and collaborators 28, 34, 30, 29. Only single-walled CNTs are considered in this work. All calculations are performed with the MESONT package 45, 46 as implemented in the LAMMPS code 47.
| entity pair | [eV] | [nm] |
| aC–aC | 0.180 | 1.327 |
| aC–CNT(16,0) | 0.386 | 1.420 |
| aC–CNT(16,0) | 0.193 | 1.420 |
| aC–CNT(32,0) | 0.440 | 2.051 |
| aC–CNT(32,0) | 0.220 | 2.051 |
Amorphous carbon particles are also represented as single beads interacting with the other entities in the system (CNTs and other aC particles) via nonbonded Lennard-Jones potentials. Their interactions are parametrized with the same procedure used for CNTs, starting from the airebo potential 48. To this end, aC particles are first modeled as a collection of carbon atoms, randomly distributed within a spherical volume, with a minimum C–C distance corresponding to a double bond (0.134 nm) (see Fig. 1). The sphere radius is set to , and the number of carbon atoms is adjusted to achieve a 2.5 gcm-3 density. The particle is then placed at varying distances from another entity, and the total interaction energy is computed as the sum of all pairwise C–C airebo contributions. This calculation is repeated 300 times with different random internal configurations of the aC structure. The resulting interaction energy profiles are averaged and fitted to determine the two Lennard-Jones parameters and used in the classical force field (see Table 1).
In this work, a single particle size and fixed density are adopted for all aC beads. Nonetheless, the same parametrization scheme can be extended to generate interaction parameters for a broad range of particle sizes and internal densities, thereby enabling flexible modeling of diverse aC morphologies relevant to different experimental conditions and device architectures.
2.3 CNT films
Film models are generated by initially placing straight CNTs and aC particles at random positions and orientations within a large simulation box, corresponding to a low-density initial configuration. The dimensions and remain fixed throughout the procedure, while only the dimension is varied during compression. The system is confined between two parallel repulsive walls in the direction, whose interactions with the particles are described by a 12-6 Lennard-Jones potentials. These walls represent the physical confinement imposed by the top and bottom electrodes. Periodic boundary conditions (PBCs) are applied only in the -plane.
To reach the target film density of either 0.3 or 0.6 gcm-3, vertical downward forces are applied to all entities in the initially expanded box while gradually moving downward the upper wall (see SI for a description of the compression protocol). The dynamics of all particles is integrated in the canonical () ensemble, with the temperature controlled by a Nosé-Hoover thermostat set at . Once the desired volume is reached, the system is equilibrated at constant temperature until the total energy fluctuates around a steady negative value, after which a final energy minimization is performed to reach a local minimum. This procedure produces a single-layer CNT film contained within a simulation box.
Multi-layer films are generated by first equilibrating four low-density layers which are then stacked together before the compression protocol (see SI), thereby yielding films with the identical overall dimensions and density that their single-layer counterparts. As discussed in section 3, structures with higher density and CNT diameter are more difficult to compact. For these cases, several compression-decompression cycles with intermediate annealing stages at elevated temperature were required to reach the desired volume.
2.4 Structural metrics
We used structural descriptors or metrics that offer a compact and interpretable representation of key morphological features and enable systematic exploration of how geometry and topology influence electron transport. The first metric quantifies the orientation of the CNTs with respect to the -plane. Because CNTs may bend and buckle, the analysis is performed on each of their segments separately. From the unitary direction vector of a segment, the angle with the -plane is calculated from its component:
where we take the absolute value to ensure that . The average and standard deviation over all segments provide a measure of the global alignment with the confinement plane and of the orientational disorder, respectively.
The second metric is the buckling factor simply defined as:
where is the number of buckling points within the -th CNT containing beads. is the total number of CNTs in the film. CNT beads are identified as buckling points when the segments they connect are misaligned by more than 28.
The third metric is the curvature factor , which quantifies the average curvature of CNTs (excluding sharp kinks). First we calculate the geometric curvature of the -th CNT, at the position of its -th bead:
where the derivative and are obtained using the finite differences method:
where is the average between the two distances and . The curvature for the first and the last beads takes the same values as the adjacent ones. The average curvature of the -th CNT is then:
where is the total number of beads in the CNT and the number of buckled beads. The average is only performed on beads that are not buckling points, since the curvature is discontinuous there and artificially large. The curvature factor for the entire film is then calculated from:
The fourth descriptor is a metric that quantifies bundling in the CNT film. The method relies on a graph-based analysis of local CNT segment connectivity. First, all CNTs are divided into segments defined by two successive beads. Then each segment is associated with its midpoint position , a unit orientation vector , and the CNT index to which it belongs. A segment is considered connected to a segment if the following conditions are met: (i) and belong to different CNTs, (ii) the distance between midpoints is less than , where is the CNT diameter (spatial proximity), (iii) the longitudinal projection is below a threshold of (the length of a segment is ) to avoid false positives from connections to the extremities of the CNTs (axial overlap), and (iv) the angle between and is smaller than the buckling angle . To prevent artificially inflated connectivity due to segment discretization, each segment is limited to at most six neighbors (the maximum number of hexagonal contacts in a bundle), and no multiple connections between the same pair of segments are allowed. Once the segment graph is built, its connected components are identified. Each component corresponds to a cluster of bundled CNT segments. The bundling metric is then defined as the average component size minus one (so that isolated segments yield zero bundling):
where is the total number of segments, the number of connected components, and the size (number of segments) of the -th component. This metric reflects the average number of connected neighbors per segment and captures local and global bundling in the film. We also report the average bundle size as an additional indicator of aggregate structure. In the limiting case where all CNT segments belong to a single connected component (i.e.: one large bundle spanning the entire system) and assuming that PBCs also apply in the -direction, the bundling metric would reach its maximum value . In the opposite case where no segment is connected to any other (i.e.: all segments are isolated) one obtains .
Finally, to characterize the electrical connectivity of the network, we analyze the ensemble of current-carrying paths (see section 2.5). Each path corresponds to a sequence of connected nodes (CNT beads) linking the two electrodes. From this set, we define two descriptors that quantify the ability of the network to conduct electricity. First, effective connectivity is a weighted average reflecting the number and quality of conductive paths. Each path contributes a weight inversely proportional to the square of the number of internal junctions (excluding the two electrode contacts), thus penalizing long and tortuous paths:
where is the number of internal junctions in the path. This descriptor favors short and direct conduction pathways. Second, minimum number of junctions corresponds to the smallest number of internal junctions observed across all conductive paths and provides a lower bound on the topological complexity required for charge transport through the network. These descriptors allow us to distinguish between highly interconnected films supporting efficient conduction and sparse networks where charge must percolate through long paths. The last descriptor is simply the smallest number of connections with either the bottom or the top electrode.
2.5 Electrical current calculation
To compute the electrical current flowing through the carbon nanotube networks, we employ a nodal analysis approach, adapted from the work of Jin et al. 49. Each CNT bead is treated as a node, and electrical connections between nodes are encoded into a sparse conductance matrix . The overall structure is modeled as a resistor network where both intratube and intertube conductances are explicitly accounted for. We define a local intratube conductance between adjacent beads of the same nanotube, and an intertube conductance between connected segments belonging to different CNTs.
Defect-free CNTs behave as one-dimensional ballistic conductors with a mean free path exceeding for the bias regime considered here 50 and a conductivity of independently of the tube length 51, 52, where is the conductance quantum. The intratube conductance is therefore given by:
where is the number of beads per CNT.
Based on the model of Durrant et al. combining density functional tight-binding (DFTB) theory and the non-equilibrium Green functions (NEGF) method 52, we estimate the average conductance at the junctions between CNTs for both chiralities as:
We note that tubes with larger diameters have lower junction conductance, an effect attributed to the reduced overlap between the delocalized conduction channels and the local contact area: as the tube circumference increases, the fraction of the electronic wave function contributing to the actual contact region decreases 52, 53.
For the junctions between CNTs and electrodes, a high conductance of 0.05 S is assigned. CNT segments are considered connected when the shortest separation between the bead-bead lines is less than a cutoff radius of nm or nm. These values correspond to twice the CNT radius plus the equilibrium van der Waals spacing in vacuum, augmented by a additional 0.15 nm. We verified that replacing this hard cutoff with an exponentially decaying distance-dependent conductance did not produce substantial differences in the total current. CNTs are also considered connected to the bottom or top electrode when their distance is less than nm. PBCs are applied in the plane of the film to ensure the correct treatment of finite-size effects, and special care is taken to correctly unwrap segment coordinates across the simulation box when identifying junctions.
Each junction is modeled by a contact conductance, which is incorporated into the matrix according to Kirchhoff’s current law. The external electrodes are represented as boundary nodes connected directly to a subset of CNTs. Once the global sparse matrix is assembled, we solve the following linear system:
where is the electrostatic potential at each node and encodes the boundary conditions imposed by the electrodes, with a bias of 0.5 V applied at the bottom electrode. The local electrical current flowing between two nodes and is given by:
As the system is modeled as a purely resistive network of Ohmic CNTs obeying Kirchhoff’s law, the total current transmitted through the film is conserved. It is therefore computed by summing all currents crossing the -plane at mid-height.
We considered two cases: (i) all CNTs are metallic, and (ii) since many CNT chiralities are semiconducting, only one third of the tubes are randomly selected as being metallic, while the remaining semiconducting ones are assumed to be non-conductive under the 0.5 V bias applied here. In the present model, the electrical conductance of each connection is treated as Ohmic and independent of the carrier statistics, such that both the fully metallic and partially metallic networks follow the same linear transport scheme. In reality, when only one third of the CNTs are metallic, the system would likely operate in a mixed regime where quantum conduction along metallic paths coexists with thermally activated hopping across semiconducting barriers 54, 55. Such effects are not explicitly captured here, but are expected to primarily rescale the overall conductance without altering the qualitative trends observed between structural descriptors and transport efficiency.
After solving the linear system given by Eq. (11), we extract the set of bonds that carry a significant current and reconstruct the connected subgraph linking the source and drain electrodes in order to identify the available dominant conduction pathways within the network. The possible contributions of aC particles to the total film conductivity are neglected in the present analysis.
3 Results and Discussion
Structures containing amorphous carbon are shown in Figs. 2 and 3 while all other configurations are provided in the SI (Figs. S3 and S4). The complete set of calculated descriptors and currents is also presented in SI (Table S1).
Due to the inherent variability in the initial random configurations and the necessity of applying a slightly different compression protocol for structures #30 and #31 (combining high-density, intermediate amorphous carbon content, and large-diameter tubes), the quantitative and physical interpretation of the descriptor correlations remains exploratory. A statistically more robust analysis would require multiple realizations per condition and a more uniform compression protocol across all structures. Nevertheless, the consistent qualitative trends observed across the dataset, in particular for buckling and bundling, indicate that the selected set of descriptors captures relevant topological and geometrical features of the CNT networks and provides a useful basis for future more systematic studies. We note that we were unable to identify the specific reason why the structures #30 and #31 developed a noticeably different morphology compared to the other aC-containing films and consider them representative structures.
Low-density films composed of short CNTs (15 or 40 nm in length; #1#4 and #17#20) tend to exhibit large voids and pronounced bundle formation, reflected in high values of the bundling factor and average bundle size , while showing no detectable buckling (). According to the stability criterion proposed by Volkov et al. 34, which identify stable networks as those that avoid fragmentation into isolated and weakly connected bundles, these systems would be considered unstable. These configurations consistently show poor effective connectivity () and low electrical conductivity, particularly for high value of . For low-density films with 100-nm long tubes we observe a reduction in bundling and a concurrent increase in bending, curvature and current for the (16,0) chirality. These trends remain nearly unchanged for the large-diameter (32,0) tubes, showing that thinner CNTs are more prone to mechanical deformation under compression, which in turn improves intertube contacts and electrical conduction.
In contrast, high-density structures exhibit significantly fewer voids, especially in the four-layer configurations and for the (32,0) chirality. This trend is not surprising. On the one hand, (32,0) films at 0.6 gcm-3 are already close to their close-packed density, whereas the corresponding value for (16,0) films is approximately 1.5 times higher, leaving more space for void formation. On the other hand, since the position of the voids within each layer is random, the voids do not align when the layers are stacked together. We note that without lateral confinement along the -direction several of these high-density films would expand and adopt a lower-density structure. If stability were judged on the basis of the ability of the films to maintain their morphology in vacuum, some high-density films would likely be considered unstable. However, in the capacitor geometry considered here, the lateral confinement is intrinsic, and the film density remains constant throughout. For these dense morphologies, layering systematically reduces electrical conductivity, whereas this trend is not observed for their low-density counterparts. The role of amorphous carbon appears to be more versatile, with its influence sensitively depending on the combined effects of chirality, mass fraction, and compression protocol. These observations highlight the complex interplay between structural parameters such as chirality, tube length, density, and amorphous carbon content that collectively dictate the final morphology and transport performance of the network.
To quantitatively assess the relationships between structural descriptors and transport properties, we first computed the Pearson correlation matrix 56 across all available parameters, as shown in Fig. 4(a). Several strong correlations emerge, providing insight into how geometry and topology affect the electrical performance of CNT networks. Local geometric distortions, captured by the curvature factor and buckling factor , are among the most influential descriptors. These two quantities are strongly correlated (), as both reflect geometric distortions at the nanotube segment level. In particular, both and show a strong positive correlation with the total current ( and , respectively) and even more with the current when only one third of the tubes are considered metallic ( and ). These results suggest that local mechanical distortions facilitate charge transport, probably by enhancing intertube overlap and contact and/or by reducing the number of junctions along the main conductive paths.
In contrast, the bundling factor shows a strong negative correlation with () and with (), indicating that highly bundled networks tend to conduct less efficiently. Similar trends have been widely reported in low-density, transparent CNT films, where increasing bundle size reduces the number of effective intertube junctions and limits charge percolation 57, 58, 59, 60, 55, 61. The geometric arguments made in these papers are consistent with the behavior we observe in our high-density films.
The effective connectivity exhibits a moderate positive correlation with both currents ( with and with ), indicating that global network connectivity contributes to transport, though to a lesser extent than local geometric factors such as curvature and buckling, which display stronger correlations. This highlights the combined influence of local and global structural features on charge conduction. Interestingly, the minimum number of connections to the electrodes shows only a weak correlation ( and , respectively), which does not rule out a more significant role of the electrodes under cycling or nonequilibrium conditions.
Global structural parameters also influence connectivity. Tube length is positively correlated with curvature (), buckling (), (), and both currents ( for and for ). Longer tubes are more likely to establish multiple contacts and undergo elastic deformation, thereby improving network connectivity. Film density correlates moderately with , , and (), as well as with total current (), consistent with the expected increase in intertube interactions in denser networks. In contrast, the number of layers is anticorrelated with connectivity and transport ( with currents, with ), likely reflecting the tendency of multilayer films to introduce additional junctions along percolation paths. Finally, the tube orientation angle and its dispersion are highly correlated with each other (), but show negligible correlation with transport properties or connectivity descriptors, indicating that the macroscopic angular disorder does not significantly affect global transport in the present dataset. However, the effects of local alignment or anisotropic percolation pathways may still play a role beyond the scope of the linear correlation analysis presented here.
The role of amorphous carbon is complex and highly structure-dependent. Although it correlates positively with curvature () and buckling (), its correlation with transport is less clear. No monotonic trend is observed between the aC content and the current. This likely reflects the inhomogeneous spatial distribution of aC and the non-uniform generation protocols. In particular, structures with chirality (32,0) and 20-30% aC content (#30 and #31) required additional pressure-compression cycles, leading to the formation of large aC clusters and modified bundle morphologies. Such artifacts likely affect transport independently of the nominal aC percentage. For (16,0) structures (#13-#16), aC consistently aligns along CNTs, whereas (32,0) structures exhibit more diverse behaviors: junction clustering at 10% aC (#29), plaque formation at 20-30% (#30-#31), and tube alignment at 40% (#32). These variabilities highlight the sensitivity of aC morphology to both diameter and processing history. Consequently, conclusions regarding the role of aC should be drawn with caution. A more rigorous assessment would require the controlled structure generation and multiple realizations per configuration to statistically capture the influence of aC on film morphology and transport. These observations also suggest that it should be possible to tune the morphology of aC-containing films through careful control of the construction protocol, enabling the modeling of different experimental scenarios depending on the targeted application or physical conditions. Finally, we note that the aC content is generally expected to reduce the conductivity of the film by introducing additional carrier scattering 62, an effect distinct from the morphological influence discussed here.
To test the robustness of the trends identified with the Pearson correlation matrix, we also calculated the Spearman rank correlation matrix 63, as shown in Fig. 4(b), which captures monotonic and possibly nonlinear relationships between variables. Overall, the Spearman analysis confirms most of the trends previously observed, particularly the strong associations between curvature, buckling, and transport properties. Some correlations involving the amorphous carbon content were slightly reduced, suggesting that its influence on transport is not strictly linear. In contrast, the correlations between angular descriptors ( and ) and current values were more pronounced by the Spearman analysis, hinting at potential nonlinear effects of tube orientation that are not well captured by Pearson correlation. This observation suggests that local alignment may influence transport pathways in a more subtle, nonlinear manner. The overall consistency between the two analyses reinforces the relevance of the identified descriptors, while also emphasizing the importance of considering nonlinear interactions in future modeling efforts.
Finally, we performed a principal component analysis (PCA) 64, 65, 66, 67, 68 on the entire set of normalized structural descriptors, including geometrical and topological quantities. Electrical currents and were not included as input variables. However, their subsequent projection into the PCA space reveals how they correlate with the underlying structural patterns. The results for the most significant principal components are shown in Table S2. This dimensionality reduction reveals dominant patterns of variation across the dataset and facilitates the interpretation of the underlying structure-property relationships. The first principal component (PC1) alone captures 38.9% of the total variance, while PC2 accounts for an additional 19.3%. Together, the first three components explain more than 71% of the variance.
Analyzing component loadings, PC1 is mainly associated with curvature, buckling, effective connectivity, and amorphous carbon content, and to a lesser extent with CNT length, density, and the number of connections to electrodes; all exhibiting negative coefficients. In contrast, features such as the bundling factor, mean bundle size, number of layers, and minimum number of connections to electrodes exhibit positive contributions. This implies that structures with large curvature and buckling, high connectivity, and minimal bundling, i.e. those with strongly negative PC1 scores, are also those showing enhanced transport performance. Negative loadings of transport-enhancing descriptors simply reflect the arbitrary orientation of PC1 in the eigen-decomposition and do not imply an inverse physical relationship (see SI). This trend aligns with our earlier correlation analysis and confirms the central role of morphological order and network connectivity in enabling efficient conduction. PC2 is largely dominated by angular disorder and, to a lesser extent, by CNT length and the number of layers. These orientation descriptors exhibit limited correlation with transport, and their prominence along PC2 reflects a distinct source of structural variation unrelated to conductivity. Higher-order components (PC3 and beyond) capture more specific effects, such as chirality, bundling size, and layer count.
In the PCA projection along the PC1-PC2 plane, shown in Fig. 5, individual structures can be grouped into clusters, in particular according to tubes length, with tightly grouped configurations observed for aC-containing structures (cluster A). Apart from structures #30 and #31, networks containing aC are the most conductive. Transport vectors (, ), when added as supplementary arrows in the correlation plot, point strongly along the negative PC1 direction, confirming that well-connected, low-bundling structures with high curvature and buckling consistently produce higher current output. In contrast, structures with weak connectivity and strong bundling, characterized by positive PC1 scores, fall into the low-performance regime. These findings highlight the value of PCA not only as a tool for reducing dimensionality but also as a means to identify collective structural signatures that govern macroscopic transport.
4 Conclusion
We have presented a mesoscopic modeling framework for simulating high-density carbon nanotube networks with different structural morphologies and amorphous carbon content and to analyze their structural characteristics and conductivity in device applications. A diverse set of 32 CNT films was constructed, systematically varying chirality, tube length, network density, and aC content. The resulting structures were characterized by a set of geometrical and topological descriptors, and electrical current through the films was calculated with a nodal analysis framework. Additionally, we investigated the role of layering by comparing single- and multi-layer films, and found that its impact on connectivity and transport depends significantly on structural context, including chirality and tube length. To account for amorphous carbon, we introduced a simple yet effective modeling strategy enabling its integration into CNT films. Our analysis reveals that aC content can significantly alter the local morphology and affect transport, although in a nonmonotonic and configuration-dependent manner.
Despite the inherent variability in initial configurations and compression protocols, our descriptor-based analysis reveals consistent trends linking morphological features to electrical transport. Curvature and buckling factors, which capture local mechanical distortions, show strong positive correlations with current, suggesting that deformation enhances intertube contact and conduction pathways. In contrast, bundling suppresses the current, most likely due to a reduction in overall connectivity across the network. Global descriptors such as CNT length and network density also play significant roles, reinforcing the idea that both local geometry and global topology jointly determine charge transport. Both Pearson and Spearman correlation matrices confirm these trends, highlighting strong monotonic and, in some cases, non-linear dependencies between structural descriptors and electrical transport. Principal component analysis further consolidates these findings, with the first component defining a structural axis characterized by high curvature, strong buckling, low bundling, and high connectivity. This reduction in dimensionality highlights that efficient transport is governed by a combination of factors rather than by any single descriptor.
In summary, our results demonstrate that high-density CNT films can be rationally designed to achieve specific target morphologies by tuning the amorphous carbon content and the compression protocol, with amorphous carbon incorporated as spherical domains of controllable size and density. We also demonstrate that a limited set of structural descriptors can effectively capture and explain the electrical behavior of such films. This study lays out a foundation for a more systematic exploration of CNT-based memristive architectures and highlights the potential of mesoscale modeling as a predictive tool for guiding nanostructure design.
Acknowledgments
We acknowledge the computational resources provided by the Aalto Science-IT project, and the use of the UCL Kathleen High Performance Computing Facility (Kathleen@UCL), and associated support services, in the completion of this work. The authors are grateful to P. A. Kloza and J. A. Elliott for valuable discussions and to A. N. Volkov and L. V. Zhigilei for introduction into CNT force fields and useful comments.
Supplementary Material for
”Mesoscopic Modeling of High–Density Carbon Nanotube Films for Memristive Device Applications”
S1 Additional figures
S2 Compression protocol
First, straight CNTs are randomly placed in the simulation volume, taking into account PBCs, followed by randomly placing aC particles in the remaining space when requested. In order to place all the entities in a reasonable amount of time, the size of the box is increased in the direction, while the lateral dimensions in the -plane are kept at 200200 nm2 during the whole process. For high-density structures, the initial thickness can be increased up to 160 nm, in which case the initial density is set to 0.075 gcm-3 (See Fig. S5). After compression, the films reached 20 nm thickness and a 0.6 gcm-3 density. For low density structures, the initial density is divided by two. A final run at room temperature followed by a energy minimization is performed in the final configuration.
To build the four-layer structures, we first place the different entities randomly in four different sublayers with an initial 20 nm thickness and a density of one-quarter of the target density. An equilibration is run independently for each layer before stacking them together and performing the compression protocol (see Figs. S6 and S7). After the whole process is finished, these four-layer structures fuse into a single-layer structure with the same dimensions and density as the structures built without layering. Nevertheless, we keep the denomination “four-layer” to differentiate the two different cases.
To facilitate the compaction of the film, vertical downward forces proportional to the mass of a particular species are applied. This is required because, if a uniform force is used on all entities, aC particles of the specific size and density we chose here tend to concentrate at the bottom of the film. We then alternate between runs with and without auxiliary forces to let the system gradually approach local minima, while the upper wall is displaced to the lower values before each run without forces. We note that simply moving the wall without additional forces often leads to numerical instabilities and takes much more time. One can reach different morphologies by changing the magnitude of these forces. Here we chose values small enough to avoid numerical instabilities and high enough to reduce the calculation time (the order of magnitude is typically between 0.2 and 0.5 eVÅ-1). It is possible to configure different force fields to reach different target structures, for instance, by adding some circular forces to the -plane to mimic experimental spin coating methods.
We note that the structures #30 and #31 got stuck in a local minimum before reaching the final dimensions. We performed some extra uncompression-equilibration-compression runs to force the structures to reach the target density. In the process, large tube bundling and aC plaques appeared, giving these two structures a morphology different from that of other aC-containing structures. We do not consider this result an artifact of this protocol and use it as an example of the variety of possible structures.
S3 Structural parameters
| 1 | 16 | 0.3 | 15 | 1 | 0 | 5344 | 14.8 | 14.9 | 0.04 | 0.00 | 0.96 | 23.9 | 0.03 | 2 | 844 | 2.1 | 753.1 |
| 2 | 16 | 0.3 | 15 | 4 | 0 | 5344 | 10.4 | 11.3 | 0.04 | 0.00 | 0.98 | 32.9 | 0.02 | 2 | 739 | 2.5 | 674.1 |
| 3 | 16 | 0.3 | 40 | 1 | 0 | 2004 | 2.1 | 2.9 | 0.32 | 0.00 | 1.00 | 143.9 | 0.01 | 8 | 948 | 0.0 | 354.1 |
| 4 | 16 | 0.3 | 40 | 4 | 0 | 2004 | 2.1 | 3.0 | 0.08 | 0.00 | 0.97 | 25.7 | 0.01 | 5 | 1647 | 0.3 | 649.6 |
| 5 | 16 | 0.3 | 100 | 1 | 0 | 802 | 2.5 | 4.0 | 0.41 | 0.29 | 0.84 | 6.1 | 0.10 | 1 | 2338 | 159.3 | 2127.5 |
| 6 | 16 | 0.3 | 100 | 4 | 0 | 804 | 2.0 | 2.4 | 0.32 | 0.07 | 0.91 | 10.4 | 0.03 | 5 | 1788 | 28.5 | 1162.9 |
| 7 | 16 | 0.6 | 15 | 1 | 0 | 10688 | 7.0 | 10.1 | 0.07 | 0.00 | 0.95 | 16.8 | 0.04 | 1 | 2919 | 21.9 | 2147.3 |
| 8 | 16 | 0.6 | 15 | 4 | 0 | 10688 | 3.7 | 5.0 | 0.09 | 0.00 | 0.98 | 42.1 | 0.01 | 6 | 1043 | 11.1 | 844.1 |
| 9 | 16 | 0.6 | 40 | 1 | 0 | 4008 | 0.9 | 2.2 | 0.07 | 0.01 | 0.99 | 53.7 | 0.01 | 5 | 4669 | 2.1 | 1539.5 |
| 10 | 16 | 0.6 | 40 | 4 | 0 | 4008 | 1.7 | 2.2 | 0.20 | 0.03 | 0.99 | 60.4 | 0.02 | 7 | 233 | 11.8 | 452.1 |
| 11 | 16 | 0.6 | 100 | 1 | 0 | 1604 | 2.4 | 4.1 | 0.46 | 0.34 | 0.84 | 6.0 | 0.11 | 1 | 5630 | 316.1 | 4357.2 |
| 12 | 16 | 0.6 | 100 | 4 | 0 | 1604 | 3.0 | 3.6 | 0.55 | 0.30 | 0.84 | 6.2 | 0.05 | 2 | 5268 | 260.3 | 3359.5 |
| 13 | 16 | 0.6 | 100 | 1 | 10 | 1443 | 4.0 | 5.6 | 0.71 | 0.69 | 0.68 | 3.1 | 0.21 | 1 | 8707 | 695.7 | 5422.6 |
| 14 | 16 | 0.6 | 100 | 1 | 20 | 1283 | 5.4 | 5.7 | 0.85 | 0.79 | 0.52 | 2.1 | 0.29 | 1 | 3064 | 836.8 | 5517.2 |
| 15 | 16 | 0.6 | 100 | 1 | 30 | 1123 | 5.4 | 5.4 | 0.90 | 0.75 | 0.41 | 1.7 | 0.29 | 1 | 1346 | 629.5 | 4207.1 |
| 16 | 16 | 0.6 | 100 | 1 | 40 | 962 | 5.9 | 6.2 | 1.00 | 1.00 | 0.25 | 1.3 | 0.53 | 1 | 3891 | 1073.3 | 5337.4 |
| 17 | 32 | 0.3 | 15 | 1 | 0 | 2672 | 13.0 | 13.0 | 0.03 | 0.00 | 0.91 | 10.4 | 0.06 | 2 | 326 | 1.7 | 210.0 |
| 18 | 32 | 0.3 | 15 | 4 | 0 | 2672 | 7.3 | 9.9 | 0.02 | 0.00 | 0.98 | 40.7 | 0.05 | 1 | 949 | 0.2 | 220.1 |
| 19 | 32 | 0.3 | 40 | 1 | 0 | 1002 | 2.9 | 4.2 | 0.04 | 0.00 | 0.93 | 13.6 | 0.19 | 1 | 564 | 84.9 | 418.7 |
| 20 | 32 | 0.3 | 40 | 4 | 0 | 1004 | 1.9 | 2.9 | 0.04 | 0.00 | 0.97 | 28.9 | 0.03 | 5 | 1149 | 0.4 | 137.3 |
| 21 | 32 | 0.3 | 100 | 1 | 0 | 401 | 0.8 | 1.8 | 0.06 | 0.00 | 0.97 | 25.5 | 0.04 | 4 | 1183 | 0.1 | 85.7 |
| 22 | 32 | 0.3 | 100 | 4 | 0 | 404 | 0.8 | 1.3 | 0.08 | 0.00 | 0.94 | 15.1 | 0.04 | 5 | 1632 | 2.1 | 135.7 |
| 23 | 32 | 0.6 | 15 | 1 | 0 | 5344 | 9.0 | 12.5 | 0.05 | 0.00 | 0.88 | 7.8 | 0.11 | 1 | 4288 | 22.8 | 991.6 |
| 24 | 32 | 0.6 | 15 | 4 | 0 | 5344 | 4.0 | 7.0 | 0.05 | 0.00 | 0.91 | 10.9 | 0.07 | 2 | 1949 | 12.1 | 636.8 |
| 25 | 32 | 0.6 | 40 | 1 | 0 | 2004 | 1.3 | 3.3 | 0.07 | 0.00 | 0.93 | 13.9 | 0.09 | 1 | 8322 | 21.7 | 667.0 |
| 26 | 32 | 0.6 | 40 | 4 | 0 | 2004 | 1.5 | 2.5 | 0.11 | 0.00 | 0.93 | 13.4 | 0.06 | 4 | 7273 | 15.9 | 617.8 |
| 27 | 32 | 0.6 | 100 | 1 | 0 | 802 | 1.3 | 2.9 | 0.17 | 0.02 | 0.93 | 13.6 | 0.14 | 1 | 1378 | 63.2 | 942.7 |
| 28 | 32 | 0.6 | 100 | 4 | 0 | 804 | 1.6 | 2.6 | 0.27 | 0.01 | 0.86 | 6.7 | 0.10 | 2 | 9246 | 37.4 | 764.7 |
| 29 | 32 | 0.6 | 100 | 1 | 10 | 722 | 4.9 | 5.2 | 0.62 | 0.12 | 0.45 | 1.8 | 0.80 | 1 | 8426 | 322.8 | 2135.5 |
| 30 | 32 | 0.6 | 100 | 1 | 20 | 642 | 0.5 | 1.1 | 0.07 | 0.00 | 0.97 | 27.7 | 0.04 | 5 | 5773 | 0.2 | 317.4 |
| 31 | 32 | 0.6 | 100 | 1 | 30 | 562 | 2.1 | 2.9 | 0.23 | 0.00 | 0.79 | 4.7 | 0.20 | 1 | 4116 | 36.9 | 600.1 |
| 32 | 32 | 0.6 | 100 | 1 | 40 | 481 | 6.1 | 5.7 | 0.67 | 0.10 | 0.16 | 1.2 | 1.00 | 1 | 3614 | 333.8 | 1740.5 |
S4 Correlation calculations
S4.1 Calculation of Pearson Correlations
The Pearson correlation coefficients 56 were computed to quantify the linear relationships between all pairs of structural descriptors and transport properties. Each descriptor was first normalized to zero mean and unit variance. The correlation matrix was then calculated using the standard Pearson formula:
where and are the values of two variables across the 32 structures, and and their respective means.
S4.2 Calculation of Spearman Rank Correlations
The Spearman rank correlation coefficients 63 were computed to capture possible non-linear but monotonic relationships between descriptors and transport quantities. For each variable, the data values were first converted into ranks, that is, replaced by their position in the ordered sequence from smallest to largest. The correlation coefficient was then evaluated as the Pearson correlation between the ranked variables:
where and denote the ranks of and .
S4.3 Principal Component Analysis (PCA)
| PC1 (38.91%) | PC2 (19.25%) | PC3 (12.97%) | PC4 (8.00%) | PC5 (6.59%) | PC6 (5.62%) | |
| 0.028 | -0.038 | 0.615 | 0.435 | -0.018 | 0.048 | |
| -0.236 | -0.142 | 0.183 | -0.402 | 0.456 | 0.320 | |
| -0.289 | -0.324 | -0.015 | 0.094 | -0.293 | -0.313 | |
| 0.207 | -0.138 | -0.016 | -0.148 | -0.513 | 0.739 | |
| aC | -0.349 | 0.006 | -0.101 | 0.323 | 0.108 | 0.181 |
| -0.017 | 0.586 | -0.061 | -0.017 | 0.026 | 0.110 | |
| 0.019 | 0.595 | 0.013 | -0.114 | 0.069 | 0.079 | |
| -0.379 | -0.053 | -0.296 | -0.064 | -0.065 | 0.053 | |
| -0.331 | 0.019 | -0.338 | -0.238 | -0.162 | -0.092 | |
| 0.400 | -0.064 | 0.084 | -0.204 | 0.007 | -0.227 | |
| 0.246 | -0.110 | -0.365 | 0.200 | 0.531 | 0.030 | |
| -0.342 | 0.057 | 0.086 | 0.387 | 0.119 | 0.264 | |
| 0.258 | -0.303 | -0.312 | 0.177 | 0.179 | 0.255 | |
| -0.195 | -0.206 | 0.355 | -0.421 | 0.258 | 0.0434 |
Principal Component Analysis 64, 65, 66, 67, 68 was used to identify the dominant modes of structural variability and their correlation with transport properties. All structural descriptors were first normalized to have zero mean and unit variance. The covariance matrix was then constructed as:
where is the data matrix containing structures and descriptors. The eigenvalue problem:
was solved to obtain the eigenvectors (principal components) and their corresponding eigenvalues which quantify the explained variance.
Each principal component (PC) corresponds to an orthogonal linear combination of the original descriptors that maximizes the variance along successive directions. PC1 captures the largest possible variance in the dataset, PC2 the second largest while remaining orthogonal to PC1, and so on. The eigenvalues quantify the proportion of total variance explained by each component, and the eigenvector loadings indicate how each descriptor contributes to that component. The total variance captured by the first principal components reflects the degree of redundancy or correlation among the structural descriptors. When several descriptors are strongly correlated, their shared variability can be summarized by a single component, leading to a higher explained variance. In contrast, if descriptors are largely independent, the variance is distributed among several components, resulting in lower individual eigenvalues.
Finally, the first two principal components (PC1 and PC2) were used to project the dataset and visualize clustering and correlations with transport quantities (see Fig. 5). The electrical currents and were not included as input variables in the PCA. Instead, they were projected a posteriori onto the (PC1, PC2) plane by computing their Pearson correlations with each principal component. This approach allows us to visualize of how transport quantities align with dominant structural patterns without influencing the decomposition itself.
We note here that the sign of the PCA loadings is arbitrary since each eigenvector can be multiplied by without altering the variance captured or the correlations between variables. Therefore, the negative coefficients of the descriptors associated with enhanced transport simply reflect the chosen orientation of PC1 in the eigen decomposition and do not imply an inverse physical relationship.
Graphical abstract